OpenStudy (anonymous):

For the given functions f and g, find the requested function. Please show all of your work.

OpenStudy (saifoo.khan):

i think i can do this one, but where are the functions?

OpenStudy (anonymous):

saiffo mind reader

OpenStudy (saifoo.khan):

LOL^

OpenStudy (anonymous):

\[f(x)=7x-6/7x-2\]

OpenStudy (anonymous):

and

OpenStudy (anonymous):

\[f(x)=\frac{7x-6}{7x-2}\]

OpenStudy (anonymous):

\[g(x)= 4x/7x-2\]

OpenStudy (saifoo.khan):

Ok, now that's the requested function?

OpenStudy (anonymous):

satellite got it

OpenStudy (saifoo.khan):

No! the question isnt complete yet!

OpenStudy (anonymous):

find f + g

OpenStudy (saifoo.khan):

Simple as cake.

OpenStudy (saifoo.khan):

\[\frac{7x-6}{7x-2} + \frac{4x}{7x-2}\] \[\frac{7x-6 + 4x}{7x-2}\]

OpenStudy (saifoo.khan):

\[\frac{11x-6}{7x-2}\]Answer

OpenStudy (anonymous):

omg wow

OpenStudy (saifoo.khan):

;)

OpenStudy (saifoo.khan):

Dont forget to hit good answer

OpenStudy (anonymous):

lol saifoo

OpenStudy (saifoo.khan):

;)